OpenStudy (anonymous):

Find the annual interest rate if a $5000 investment earns $550 annual simple interest in one year. (Points : 1) 10.5% 11% 11.5% 12%

OpenStudy (anonymous):

11%

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

150/100=1.5 so 1.5 leaves is equal to 1% 177-150=27 her collection grew by 27 leaves 27/1.5=percent

OpenStudy (anonymous):

Any more?